# Fairness is Not Silence: Unmasking Vacuous Neutrality in Small Language Models

### v1

@@ Abstract
The rapid adoption of Small Language Models (SLMs) for on-device and resource-constrained deployments has outpaced our understanding of their ethical risks. To the best of our knowledge, we present the first large-scale audit of instruction-tuned SLMs spanning 0.5 to 5 billion parameters, an overlooked “middle tier” between BERT-class encoders and flagship LLMs. Our evaluation includes nine open-source models from the Qwen 2.5, LLaMA 3.2, Gemma 3, and Phi families. Using the BBQ benchmark under zero-shot prompting, we analyze both utility and fairness across ambiguous and disambiguated contexts. This evaluation reveals three key insights. First, competence and fairness need not be antagonistic: Phi models achieve $\geq$ 90% F1 scores while exhibiting minimal bias, showing that efficient and ethical NLP is attainable. Second, social bias varies significantly by architecture: Qwen 2.5 models may appear fair, but this often reflects vacuous neutrality, random guessing or evasive behavior, rather than genuine ethical alignment. In contrast, LLaMA 3.2 models exhibit stronger stereotypical bias, suggesting overconfidence rather than neutrality. Third, compression introduces nuanced trade-offs: 4-bit AWQ quantization improves F1 scores in ambiguous settings for LLaMA 3.2-3B, but increases disability-related bias in Phi-4-Mini by over 7 percentage points. These insights provide practical guidance for the responsible deployment of SLMs in applications demanding fairness and efficiency, particularly benefiting small enterprises and resource-constrained environments.
Fairness is Not Silence: Unmasking Vacuous Neutrality in
@@
Sumanth Manduru  Carlotta Domeniconi George Mason University {smanduru, cdomenic}@gmu.edu
@@ 1 Introduction
Large Language Models (LLMs) have achieved impressive performance across a wide range of natural language processing (NLP) tasks, including question answering (QA) Grattafiori et al. (2024); OpenAI et al. (2024). These models are trained using self-supervised learning on vast amounts of unlabelled data, allowing them to effectively learn language patterns through methods like masked language modeling Devlin et al. (2019a). However, as LLMs increase in size, they become more prone to inheriting social biases from the training data Guo et al. (2024). These biases may manifest when LLMs respond to questions involving socially sensitive content, potentially leading to biased and harmful outputs Kaneko and Bollegala (2021); Delobelle and Berendt (2022). These risks are especially concerning in high-stakes applications like medical diagnostics Schmidgall et al. (2024), where maintaining fairness and robustness is critical Liang et al. (2023).
Despite their powerful capabilities, LLMs face challenges when deployed locally due to high computational demands Chien et al. (2023); Zhu et al. (2024). To address this issue, researchers have shifted focus towards developing smaller, more efficient models called Small Language Models (SLMs). These efficient models are often the result of a multi-stage process that involves pre-training and compressed versions of larger models Llama3.2 (2024); GemmaTeam et al. (2025), or can be trained directly as compact networks Abdin et al. (2024a, b); Qwen et al. (2025). Their fast inference and low resource requirements make them well-suited for deployment on edge devices such as smartphones and embedded systems. Recent light weight models, like LLaMA3.2-1B and 3B, offer features such as multilingual generation, tool integration, and autonomous agent-like behavior while also significantly reducing environmental impact.
Many SLMs are developed through model compression techniques aimed at reducing size and computational requirements while preserving performance. Methods such as pruning, knowledge distillation, and quantization are frequently utilized for this purpose. Pruning methods, like Wanda Sun et al. (2024) and SparseGPT Frantar and Alistarh (2023) efficiently reduce model parameters while maintaining accuracy, whereas quantization techniques like AWQ Lin et al. (2024a) decrease memory usage by lowering bit precision during inference. Knowledge distillation Hinton et al. (2015) involves training a smaller model to replicate the performance of a larger, pre-trained model. However, these compression techniques can unintentionally influence model fairness, highlighting the need for rigorous assessments of both performance and social bias in SLMs Gonçalves and Strubell (2023).
Although research on bias and fairness has rapidly advanced, most studies have focused on large models (8B parameters and above) Huang et al. (2023); Hong et al. (2024); Gallegos et al. (2024b), or smaller models such as BERT (typically under 0.5B parameters) Parrish et al. (2022); Gonçalves and Strubell (2023), leaving a gap for intermediate-sized models. These lightweight models, typically ranging from 0.5B to 5B parameters, are gaining importance for practical applications as they strike a balance between computational efficiency and robust language processing. Given their potential for real-world deployment, particularly within small and medium enterprises (SMEs), it is essential to thoroughly evaluate their robustness and fairness. Our main contributions and observations are summarized as follows:
  • We demonstrate that competence and fairness can be mutually inclusive. Phi models achieve strong performance with minimal bias, demonstrating that ethical and effective NLP is feasible even under ambiguity.
  • We uncover the phenomenon of vacuous neutrality, where models like Qwen appear fair, consistently scoring near-zero bias under both ambiguous and disambiguated conditions—but do so by relying on conservative or random responses. This behavior sacrifices specificity and usefulness, revealing a gap between perceived neutrality and meaningful fairness.
  • We reveal significant architecture-dependent biases. LLaMA3.2-3B and Qwen2.5-3B struggle to interpret bias-related uncertainty, leading to stronger stereotypical responses. In contrast, Phi-4-Mini shows greater stability and fairness across demographics, effectively handling ambiguity and making it better suited for fairness-critical applications.
  • We observe nuanced compression trade-offs: 4-bit AWQ quantization affects utility and fairness unevenly across models. Phi-4-Mini suffers from performance degradation and variable fairness outcomes, while LLaMA3.2-3B retains utility in ambiguous settings and exhibits reduced bias. This underscores the need for fairness-aware evaluation when compressing SLMs.
@@ 2 Related Work
Social Bias in LLMs
@@
Numerous studies have shown that LLMs not only reflect existing social biases in their responses, particularly around sensitive attributes such as gender, race, and sexual orientation, but can also amplify these biases during downstream tasks Venkit et al. (2023); Gonçalves and Strubell (2023). Multiple evaluation frameworks were introduced to address this issue such as StereoSet Nadeem et al. (2020) and UNQOVER Li et al. (2020). These studies analyzed prominent transformer-based language models, such as BERT Devlin et al. (2019b), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), and GPT-4 Törnberg (2023), revealing varying levels of social bias within these models. The findings indicate that, despite architectural advancements, notable biases persist. Moreover, this evaluation demonstrated that even models subjected to fine-tuning and filtering can still harbor social biases.
Impact of Model Compression on Social Bias
@@
Model compression techniques can have unintended consequences for fairness measures. Some studies have shown that compression strategies may exacerbate social biases in language models Ramesh et al. (2023) and cause unpredictable shifts in model behavior Xu et al. (2024). However, other research suggests that compression can also act as a regularizer, potentially reducing bias in certain self-supervised models. For example, Lin et al. (2024b) reveal that by using methods such as row-pruning and training wider, shallow models can effectively mitigate social bias within self-supervised learning (SSL) frameworks. This duality arises because compression techniques can either act as a regularizer, reducing overfitting and thus mitigating bias, or distort model representations, inadvertently amplifying existing biases. Therefore, the effect of compression on social bias is inherently complex and context-dependent.
While numerous studies Gallegos et al. (2024a); Li et al. (2023) have confirmed the presence of social bias within LLMs, how compression techniques affect bias, either by exacerbating or mitigating it, in SLMs of the proposed sizes remains relatively underexplored. Most existing research in this domain has focused on either compressing very large models (8B parameters and above) Hong et al. (2024) or evaluating smaller models like BERT (less than 0.5B parameters) Gonçalves and Strubell (2023), leaving a significant gap in understanding the intermediate range. To bridge this gap, we aim to systematically evaluate open-source light-weight models ranging from 0.5B to 5B parameters, with a focus on examining how these models exhibit social bias.
@@ 3 Empirical Evaluation
In our experiments we investigate the following research questions regarding the fairness and utility of SLMs under realistic deployment constraints:
RQ1: How do lightweight instruction-tuned language models (0.5B–5B) perform in terms of task competence and social bias, particularly in ambiguous reasoning scenarios?
RQ2: How are the predictions of these models distributed across the categories of social bias and answer choices on the QA tasks with zero shots?
RQ3: What are the effects of model compression, specifically 4-bit AWQ quantization, on both utility and fairness across different model families?
@@ 3.1 Language Models (LMs)
We evaluate a diverse set of nine instruction-tuned language models (LMs) from four prominent families: Qwen2.5, LLaMA3.2, Gemma3, and Phi. These models span a range of sizes and architectures, allowing us to systematically investigate how social bias manifests across different parameter scales. For structured comparison, we categorize the models into two tiers: Tiny models (0.5B–2B parameters), including Qwen2.5-0.5B, Qwen2.5-1.5B, Gemma3-1B, and LLaMA3.2-1B; and Small models (2B–4B parameters), including Qwen2.5-3B, Gemma3-4B, LLaMA3.2-3B, Phi-3.5-Mini, and Phi-4-Mini. All models are evaluated under zero-shot prompting conditions, without any task-specific fine-tuning. To ensure robustness, each evaluation is repeated across 10 randomized trials, where samples from each demographic category are independently shuffled in every run.
3.2 Dataset
In this study, we use the BBQ dataset Parrish et al. (2022), a critical multiclass benchmark for evaluating social biases exhibited by LMs in QA tasks Xu et al. (2024); Liang et al. (2023). BBQ is particularly valuable because it reflects real-world scenarios in which demographic cues may be either implicit or explicitly stated. The BBQ dataset comprises natural language questions spanning 11 distinct demographic categories, including two intersectional categories: Race × Gender and Race × Socioeconomic status (SES). Each question in the dataset is provided in two distinct contexts: an Ambiguous Context, in which demographic information is implied implicitly, and a Disambiguated Context, where demographic details are explicitly specified. Each question contains three candidate answers: (1) a bias-reinforcing answer (Target), (2) a bias-negating answer (Non-Target), and (3) an "Unknown" option, indicating ambiguity. The positions of these candidate answers are randomized within the dataset to prevent positional bias during evaluation.
@@ 3.3 Evaluation Metrics
All the evaluation metrics follow the definitions provided in the BBQ benchmark by Parrish et al. (2022). This ensures consistent and comparable measurement of model fairness and task competence under both ambiguous and disambiguated contexts. Task competence is measured using the F1 score, which is the harmonic mean of precision and recall, computed across categories and contexts. Fairness is quantified using a Bias Score, which captures the extent to which a model exhibits stereotypical behavior. A positive score indicates alignment with social stereotypes, while a negative score reflects an inverse or anti-stereotypical tendency. In disambiguated contexts, the bias score is computed as:
@@
See Appendix Section A for dataset and evaluation details. Figure  5 presents example prompts from the BBQ dataset. The code is available at: https://anonymous.4open.science/r/BBQ-on-SLMs-E6E0.
4 How do Tiny and Small Language Models behave with respect to social bias when evaluated on the BBQ dataset?
Figure 1: Bias scores for (a) Tiny LMs (the first two heatmaps) and (b) Small LMs (the last two heatmaps). Rows corresponds to a social bias category, and columns to instruction-tuned models. The heatmaps reflect bias scores under Ambiguous context (first and third) and Disambiguated context (second and fourth). Red shades indicate stereotypical alignment, blue denotes anti-stereotypical responses, and pale or gray cells represent near-neutral outputs. Most scores lie within ±15%.
In Figure 1, we observe that the Qwen family (e.g., Qwen2.5-1.5B) exhibits the lowest measurable bias, with near-zero scores across all categories in the ambiguous context, indicating no detectable bias at the resolution of this evaluation. This neutrality suggests that Qwen models may adopt a generalized or non-committal response strategy when demographic details are present, minimizing harm but potentially reducing expressiveness or informativeness. The Phi series ranks next in bias control, maintaining relatively low and balanced scores across most dimensions. In contrast, the LLaMA family displays the most pronounced biases among both Tiny and Small models, with consistent stereotyping across multiple social categories, regardless of context. The Gemma models, although slightly less biased than LLaMA, still exhibit substantial stereotypical alignment, particularly in their smaller variants.
When comparing ambiguous and disambiguated contexts, we observe that disambiguation often amplifies bias scores in both stereotypical and counter-stereotypical directions. For instance, in the Religion category, Phi-series models exhibit minimal bias when religious identity is ambiguous but show increased stereotypical responses when a specific affiliation is stated. This suggests that explicit demographic cues can inadvertently trigger bias-aligned behavior, underscoring the importance of careful prompt design. Similarly, in categories like Disability Status, SLMs such as LLaMA and Gemma occasionally demonstrate anti-stereotypical behavior, indicating that contextual clarity may enable larger models to leverage counter-stereotypical reasoning.
When analyzing bias by category, we find that some social categories consistently generate strong bias responses, while others remain relatively neutral across models. Physical appearance emerges as the most bias-sensitive category, for instance, Gemma3-1B records the highest bias scores, with +12.2% in ambiguous and +14.4% in disambiguated contexts. These results indicate that models tend to align with stereotypes when encountering descriptors related to body weight, visible disabilities, or non-normative traits (e.g., short stature, strabismus), reflecting internalized cultural associations. In the Disability Status category, we observe a surprising trend: bias behavior varies significantly with model size. Smaller LLaMA and Gemma models tend to reinforce stereotypes, while their larger counterparts exhibit anti-stereotypical behavior, resulting in negative bias scores. This suggests that larger models may have developed stronger ethical safeguards, possibly due to additional data or refined instruction tuning. Other categories, such as Age, SES, Gender Identity, and Nationality, show moderate but consistent bias, making them important to monitor in sensitive applications. In contrast, categories related to Race and Sexual Orientation consistently yield low bias scores, even under disambiguation. Whether this neutrality stems from balanced training data, or effective alignment remains unclear, but the consistent absence of bias across models is a promising outcome.
4.1 How competent are Tiny and Small Language Models in reasoning under social bias scenarios?
The goal is to assess whether low-bias models exhibit sufficient task competence or if fairness metrics obscure random or suboptimal behavior. Although small models typically outperform tiny models, our F1 heatmaps in Figure 2 show that increased model size alone does not guarantee higher competence, especially under ambiguity. Tiny models often score around 15–17% F1, indicating near or below random guessing. However, disambiguated contexts significantly boost performance, with tiny models reaching around 40% F1 and small models achieving between 80–95%. For instance, LLaMA3.2-3B scores below 9% F1 on ambiguous "Age" and "Nationality" tasks but surpasses 80% with clear context, demonstrating sensitivity to explicit input.
The Gemma3 family displays an intriguing reversal in performance trends. Under ambiguous conditions, Gemma3-1B surprisingly outperforms its larger 4B counterpart. For example, on “SES,” the 1B model scores 37.2% F1, compared to just 11.7% for the 4B variant. However, with disambiguated inputs, Gemma3-4B significantly improves, averaging 87.5% F1, a substantial increase over both its ambiguous-context performance and the 1B variant. This pattern suggests the smaller model might rely on subtle cues or training artifacts, while the larger model performs best with explicit context.
In contrast, the Qwen2.5 family consistently underperforms, scoring around 16% F1 across ambiguous settings and only marginally higher ( 20%) when contexts are disambiguated. In particular, scaling does not enhance performance. Qwen2.5-1.5B even regresses relative to the 0.5B variant in some categories. This stagnation suggests that Qwen models may be undertrained for bias-sensitive reasoning tasks, lacking both robust ambiguity handling and effective context utilization.
Conversely, the Phi family demonstrates that high fairness and strong task competence can coexist. Phi-3.5-mini frequently achieves over 90% F1 in ambiguous contexts such as “Gender Identity,” “Nationality,” “Race × Gender,” and “Sexual Orientation,” while Phi-4-mini similarly excels in disambiguated settings (e.g., 98.8% on “SES”). Nevertheless, both Phi models consistently underperform by 10–15% in the “Physical Appearance” category, suggesting residual stereotypical biases despite overall robustness.
These findings notably highlight that bias behavior and model utility does not correlate directly with model size. For example, LLaMA3.2-3B-Instruct, compressed via pruning and distillation, then aligned using Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and safety tuning, exhibits stronger stereotypical alignment and low performance. In contrast, Phi-4-Mini, trained from scratch with SFT, Direct Preference Optimization (DPO), and rigorous safety measures, demonstrates more fairness and competence. This underscores the greater influence of alignment strategy and architecture over parameter count alone.

### v2

# Beyond Bias Scores: Unmasking Vacuous Neutrality in Small Language Models

@@ Abstract
The rapid adoption of Small Language Models (SLMs) for resource constrained applications has outpaced our understanding of their ethical and fairness implications. To address this gap, we introduce the Vacuous Neutrality Framework (VaNeu), a multi-dimensional evaluation paradigm designed to assess SLM fairness prior to deployment. The framework examines model robustness across four stages - biases, utility, ambiguity handling, and positional bias over diverse social bias categories. To the best of our knowledge, this work presents the first large-scale audit of SLMs in the 0.5–5B parameter range, an overlooked “middle tier” between BERT-class encoders and flagship LLMs. We evaluate nine widely used SLMs spanning four model families under both ambiguous and disambiguated contexts. Our findings show that models demonstrating low bias in early stages often fail subsequent evaluations, revealing hidden vulnerabilities and unreliable reasoning. These results underscore the need for a more comprehensive understanding of fairness and reliability in SLMs, and position the proposed framework as a principled tool for responsible deployment in socially sensitive settings. The code is available at: https://anonymous.4open.science/r/Vacuous-Neutrality-Framework-5314/.
Beyond Bias Scores: Unmasking Vacuous Neutrality in
@@
Sumanth Manduru George Mason University smanduru@gmu.edu          Carlotta Domeniconi George Mason University cdomenic@gmu.edu
@@ 1 Introduction
Large Language Models (LLMs) have achieved state-of-the-art performance across a wide range of natural language processing tasks, from question answering (QA) to multilingual generation grattafiori2024llama3herdmodels; openai2024gpt4technicalreport. Trained on massive unlabelled corpora, these models excel at capturing linguistic patterns through self-supervised learning objectives such as masked language modeling devlin-etal-2019-bert. However, their scale brings two major challenges. First, LLMs are computationally expensive to deploy locally, limiting accessibility 10.1145/3604930.3605705; zhu2024surveymodelcompressionlarge. Second, their reliance on large-scale web data makes them prone to reproducing and amplifying harmful social biases, with fairness risks in high-stakes settings such as healthcare and education kaneko2021unmaskingmaskevaluating; schmidgall2024addressingcognitivebiasmedical.
To overcome the computational barrier, researchers have increasingly turned to SLMs typically under 5B parameters that offer faster inference, lower memory requirements, and reduced environmental impact. SLMs emerge either through compressing larger LLMs meta_llama_3_2_connect; gemmateam2025gemma3technicalreport, or by training compact architectures from scratch abdin2024phi3technicalreporthighly; qwen2025qwen25technicalreport. Their efficiency makes them particularly attractive for deployment on edge devices, where resources are constrained but fairness and robustness remain critical. Most SLMs rely on compression techniques such as pruning, quantization, and knowledge distillation to balance efficiency with accuracy. Yet, compression is not fairness-neutral: pruning strategies like Wanda sun2024simpleeffectivepruningapproach or SparseGPT frantar2023sparsegptmassivelanguagemodels, and quantization methods like AWQ lin2024awqactivationawareweightquantization, may inadvertently reshape model biases. This highlights the need to jointly assess performance and fairness in SLMs rather than privileging only one direction goncalves-strubell-2023-understanding.
Figure 1: The Vacuous Neutrality Framework (VaNeu): a four-stage evaluation paradigm for assessing SLMs across Bias, Utility, Ambiguity Handling, and Positional Bias. Stage 1 (Bias) examines fairness via bias score, Stage 2 (Utility) tests task competence using F1 score, Stage 3 (Ambiguity Handling) measures calibrated caution via Target-to-NonTarget Ratio (TNR) and Unknown Ratio (UR), and Stage 4 (Positional Bias) evaluates response distribution consistency using normalized KL divergence.
While bias and fairness evaluations have been extensively conducted on very large models (8B+) huang2023trustgptbenchmarktrustworthyresponsible; gallegos2024selfdebiasinglargelanguagemodels and smaller models under 0.5B parameters such as BERT parrish-etal-2022-bbq, the intermediate range of 0.5B–5B remains largely understudied-despite its growing significance for practical deployment. These mid-sized models strike a balance between efficiency and capability, making them especially relevant for real-world applications. This gap raises an important question: Can these SLMs be trusted in socially sensitive settings?
To address this, we introduce an evaluation paradigm, referred to as the Vacuous Neutrality Framework (VaNeu), that jointly examines bias, utility, ambiguity handling, and positional bias. Applying this framework enables a more scrutinized assessment of SLMs and provides insights into whether they can be reliably deployed without sacrificing fairness and ethical considerations. Our main contributions are summarized as follows:
$\bullet$ We introduce the Vacuous Neutrality Framework (VaNeu), a multi-stage evaluation approach that assesses SLMs across four key dimensions: Bias, Utility, Ambiguity handling, and Positional bias.
$\bullet$ We conduct a systematic evaluation of nine mid-sized transformer-based SLMs (0.5B–5B), an underexplored but increasingly important class of models for practical deployment, using socially sensitive benchmarks (e.g., BBQ, StereoSet, and CrowS-Pairs).
$\bullet$ We identify critical trade-offs across SLMs. In some cases, models demonstrate high task performance with minimal bias, suggesting that competence and fairness can align even under ambiguity. In other cases, models register bias scores close to zero but exhibit vacuous neutrality, appearing unbiased through conservative or random predictions, which reduces specificity and usefulness.
More broadly, Our analysis highlights variation across model families, sizes, and datasets, underscoring that fairness behaviors are not uniform among these SLMs. These findings provide guidance for the responsible use of SLMs in socially sensitive applications.
@@ 2 Related Work
Social Bias in LLMs:
@@
Numerous studies have shown that LLMs not only reflect existing social biases in their responses, particularly around sensitive attributes such as gender, race, and sexual orientation but can also amplify these biases during downstream tasks venkit2023nationalitybiastextgeneration; goncalves-strubell-2023-understanding. To evaluate such risks, several benchmarks have been developed, including StereoSet nadeem2020stereosetmeasuringstereotypicalbias and UNQOVER li-etal-2020-unqovering. Analyses of prominent transformer-based models such as BERT devlin2019bertpretrainingdeepbidirectional, RoBERTa liu2019robertarobustlyoptimizedbert, GPT-2 Radford2019LanguageMA, and GPT-4 törnberg2023chatgpt4outperformsexpertscrowd reveal that, despite architectural advancements and mitigation strategies such as fine-tuning or data filtering, notable biases persist. These findings highlight that fairness challenges remain deeply embedded across model families and scales.
Impact of Model Compression on Social Bias:
@@
Model compression techniques, while essential for improving efficiency, can have unintended consequences for fairness. Some studies show that compression strategies exacerbate social biases in language models ramesh-etal-2023-comparative and cause unpredictable shifts in behavior xu2024perplexitymultidimensionalsafetyevaluation, whereas others suggest compression may act as a regularizer, mitigating bias in certain contexts Lin_2024. This duality arises because compression can either reduce overfitting and thereby dampen bias, or distort learned representations in ways that amplify it. Thus, the fairness implications of compression are complex and highly context-dependent.
While numerous studies confirm the persistence of social bias in LLMs gallegos2024biasfairnesslargelanguage; Li2023ASO, relatively little is known about how these biases manifest in SLMs. Existing work has predominantly focused on large-scale models (8B+ parameters) hong2024decodingcompressedtrustscrutinizing or on much smaller models such as BERT (under 0.5B parameters) goncalves-strubell-2023-understanding. This leaves a significant gap in understanding mid-sized SLMs (0.5B–5B), a model class that is increasingly attractive for deployment due to its balance of efficiency and capability. To address this gap, we conduct a systematic evaluation of open-source, transformer-based SLMs within this intermediate range, focusing specifically on their tendencies to exhibit social bias under socially sensitive benchmarks. To the best of our knowledge, this is the first comparative fairness audit spanning multiple transformer families of SLMs in the 0.5B–5B parameter range across widely used bias evaluation benchmarks.
3 Vacuous Neutrality Framework
Evaluating SLMs requires going beyond single dimension metrics. We introduce the Vacuous Neutrality Framework (VaNeu), a multi-stage evaluation paradigm designed to assess SLMs across four complementary dimensions: Bias, Utility, Ambiguity Handling, and Positional Bias.
3.1 Bias
The first dimension, bias, examines whether a model disproportionately favors stereotypical completions over anti-stereotypical or neutral alternatives. Such behavior suggests reliance on social associations encoded in training data rather than task-relevant reasoning. Bias is particularly concerning because it often arises in sensitive categories such as gender, race, religion, sexual orientation, and socioeconomic status. If left unaddressed, these disparities can lead not only to overtly harmful outputs but also to subtle distortions in downstream tasks such as question answering. In our framework, bias metrics are calculated to quantify this behavior, allowing us to assess whether SLMs risk reinforcing harmful stereotypes or can instead provide more balanced and fair predictions in socially sensitive contexts.
3.2 Utility
After assessing bias, we turn to the question of competence. The utility dimension evaluates whether a model can successfully accomplish its intended task. It reflects the accuracy and reliability of outputs when tested on benchmark datasets. While bias highlights disparities across sensitive categories, utility emphasizes overall effectiveness whether the system interprets inputs correctly and generates responses aligned with ground truth. Strong utility is essential for deployment, since a model that appears fair but lacks competence offers limited real-world value. In our framework, utility metrics quantify task performance, ensuring that fairness assessments are interpreted in the context of verified task competence.
3.3 Ambiguity Handling
The third dimension, Ambiguity Handling, examines how models respond to underspecified inputs. This dimension captures whether a model can recognize when “Unknown” is the appropriate answer, rather than overcommitting to a potentially biased choice or defaulting toward stereotype versus anti-stereotype options. At the same time, models should still make specific predictions when sufficient context is available. To quantify this, we assess ambiguity handling by measuring how often models abstain with ‘Unknown’ in ambiguous contexts and how reliably they prefer the intended target over non-target options when the answer is clear. Together, these measures reveal whether a model balances caution with specificity, providing insight into its robustness under uncertainty.
3.4 Positional Bias
The fourth dimension in the framework is Positional Bias. In multiple-choice settings, models may show a tendency to prefer certain answer positions (e.g., consistently selecting option “A”) while neglecting others, leading to skewed rather than balanced distributions. Such skew suggests reliance on superficial heuristics rather than genuine reasoning. Beyond affecting performance, positional bias indicates a model’s adherence to instructions. We measure this by comparing the distribution of predictions across answer positions {A, B, C} against expected baselines. This analysis highlights whether models distribute attention appropriately or rely on positional shortcuts, providing insight into both robustness and instruction-following capability.
Each dimension in this task-agnostic and dataset-agnostic framework captures a distinct aspect of model behavior, and together they offer a holistic perspective on whether SLMs can be deployed responsibly in socially sensitive applications.
Figure 2: Heatmaps show bias scores for (a) Tiny and (b) Small LMs under Ambiguous and Disambiguated contexts. Rows denote social bias categories and columns denote SLMs. Red indicates stereotypical, blue anti-stereotypical, and gray near-neutral responses. Most scores fall within ±15%, with the range spanning –100% to +100%.
@@ 4 Empirical Evaluation
In our experiments we investigate the two research questions (RQs) regarding the fairness and task competence of SLMs under realistic deployment constraints:
RQ1: How do SLMs (0.5B–5B) behave across the dimensions of the VaNeu - Bias, Utility, Ambiguity Handling, and Positional Bias?
RQ2: Are these fairness behaviors consistent across bias categories, model families, and parameter scales or do they vary in systematic ways?
@@ 4.1 Language Models (LMs)
We evaluate a diverse set of nine instruction-tuned SLMs from four prominent families: Qwen2.5, LLaMA3.2, Gemma3, and Phi. These models span a range of sizes and families, allowing us to systematically investigate how social bias manifests across parameter scales. For structured comparison, we categorize the models into two tiers: Tiny models (0.5B–2B parameters), including Qwen2.5-0.5B, Qwen2.5-1.5B, Gemma3-1B, and LLaMA3.2-1B; and Small models (2B–4B parameters), including Qwen2.5-3B, Gemma3-4B, LLaMA3.2-3B, Phi-3.5-Mini, and Phi-4-Mini. All models are evaluated in a zero-shot multiple-choice format using consistent prompts across datasets, without any task-specific fine-tuning. Decoding is performed with greedy search (temperature = 0.0, top-p = 1.0) to ensure reproducibility and eliminate sampling variance. To ensure robustness, each evaluation is repeated across 10 randomized trials, where samples from each demographic category are independently shuffled in every run.
4.2 Datasets
We evaluate models on three socially sensitive benchmarks that differ in task structure and ground truth, but are cast into a unified multiple-choice QA format for consistency across SLMs.
BBQ (Bias Benchmark for QA) parrish-etal-2022-bbq: A large-scale question answering dataset designed to test stereotypical reasoning under both ambiguous and disambiguated contexts. Each instance pairs a question with demographic attributes such as gender, race, religion, or nationality. Ground truth labels are provided at the question level, which enables direct evaluation of both bias (e.g., Bias Score) and utility (e.g., Accuracy and F1 Score). BBQ is also the only dataset among the three that natively supports ambiguity handling, since it includes cases where the correct answer is “Unknown.”
StereoSet nadeem2020stereosetmeasuringstereotypicalbias: A benchmark for measuring stereotypical bias in natural language understanding. Each context is paired with candidate completions that may be stereotypical, anti-stereotypical, or unrelated. Ground truth is provided only at the level of stereotypicality, that is, whether a completion reflects a stereotype, an anti-stereotype, or an unrelated association, rather than specifying a task-correct answer. This structure makes StereoSet well-suited for evaluating bias tendencies, but less informative for measuring utility or ambiguity handling without modification.
CrowS-Pairs nangia2020crows: A minimal-pair dataset where each instance contrasts a biased and an unbiased alternative differing only by a single lexical substitution. Ground truth is provided only at the level of stereotype polarity, whether a sentence is stereo or anti-stereo, rather than specifying a task-correct answer. This design enables precise bias quantification, but does not natively support evaluation of utility or ambiguity handling.
@@ 4.3 Evaluation Metrics
We evaluate SLMs across the four dimensions of the VaNeu Framework. Each dimension is measured using benchmark-defined metrics where available (e.g., Bias Score in BBQ) and established evaluation practices to capture model behavior comprehensively. Below, we provide the equations and definitions, grouped by framework dimension.
Bias Dimension All bias metrics follow the definitions provided by the respective benchmarks. For StereoSet and CrowS-Pairs, we adopt the benchmark-defined Stereo Score, which ranges from 0 to 1, a score of 0.5 indicates neutrality, values above 0.5 indicate a preference for stereotypical completions, and values below 0.5 indicate a preference for anti-stereotypical completions. For BBQ, we use the benchmark-defined Bias Score, which ranges from –100% to 100%. Positive values indicate alignment with social stereotypes, while negative values indicate an anti-stereotypical tendency. In disambiguated contexts, the bias score is computed as: